A pendulum is hanging from a ceiling of a car having an acceleration a0 with respect to the road. Find the angle made by the string with the vertical. 

Grade:12

1 Answers

Arun
5 years ago
Dear student
 
tan theta = F/ mg
 
tan theta= mao / mg = ao /g
 
hence theta = tan-1 (ao /g)
 
Hope it helps
 
Regards
Arun (askIITians forum expert)